OpenStudy (iconforhire56):

Billy has 1 gallon of paint. He is going to pour it into a paint tray that measures 10 inches wide, 14 inches long, and 4 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen? Which of the following scenarios will happen? Billy has 1 gallon of paint. He is going to pour it into a paint tray that measures 10 inches wide, 14 inches long, and 4 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Which of the following scenarios will happen?

OpenStudy (iconforhire56):

Which of the following scenarios will happen? The paint will not fill the tray by 441 cm3. The paint will not fill the tray by 10.53 in3. The paint will fill the tray exactly. The paint will overfill the tray by 10.53 in3.

OpenStudy (nick56x):

Okay! Bear with me as I try to explain this as simply as possible!

OpenStudy (nick56x):

So, in essence, a tray is a 3D object that is a rectangular prism. And the question is asking for volume to determine which scenario will or will not happen. The volume of a rectangular prism is V = lwh . The first thing that we need to do is take a look at our dimensions: Width: 10 inches Length: 14 inches Depth: 4 centimeters The first thing that you (should) notice is that all of the units are not in the same system. SO! We need to convert our depth into inches, which is quite simple. 1 centimeter = 0.39370. 4 * 0.39370 = 1.5748 inches. The hard part is over! Now that everything is in the same system, just plug it into V = lwh: V = 10(14)(1.5748) = 220.47 inches cubed. Since 1 gallon is equal to 231 inches cubed, to figure out which scenario will happen, simply subtract your volume from the 1 gallon: 231 - 220.47 = 10.53 inches cubed. Your answer choice will be B.) The paint will not fill the tray by 10.53 in3. Hope this helps! :)